# {adjustwidth}0cm-0.2cm Higgs-Induced Gravitational Waves: the Interplay of Non-Minimal Couplings, Kination and Top Quark Mass

### v1

@@ 5 Stability beyond the Standard Model
The non-minimal coupling of the Higgs field to gravity makes the model under consideration intrinsically non-renormalisable, requiring it to be treated as an effective field theory valid up to a specific cut-off scale, $\Lambda_{\xi}=M_{P}/\xi$ Barbon:2009ya ; Burgess:2009ea ; Hertzberg:2010dc ; Burgess:2010zq ; Bezrukov:2010jz ; Bezrukov:2012hx ; Ren:2014sya ; Bezrukov:2014ipa ; Fumagalli:2017cdo ; Antoniadis:2021axu ; Mikura:2021clt ; Ito:2021ssc ; Escriva:2016cwl ; Karananas:2022byw , which may either signal the transition to a strongly coupled regime Aydemir:2012nz ; Calmet:2013hia ; Escriva:2016cwl ; Saltas:2015vsc or the emergence of new degrees of freedom beyond the SM content Giudice:2010ka (for a review in the context of Higgs inflation see Rubio:2018ogq ). For the non-minimal couplings considered in this work, $\Lambda_{\xi}$ is estimated to range from approximately $10^{16}$ to $10^{17}$ GeV, i.e. typically well above the SM vacuum instability scale. However, the validity of the low-energy effective field theory could be further reduced in the presence of additional sub-Planckian degrees of freedom beyond the SM content.
@@
Figure 6: Impact of non-renormalisable sextic operators in the vacuum structure of the Higgs effective potential. The parameter space above the black line allows for the presence of only one minimum at high scales. The red area indicates that the instability scale is probed already during the inflationary stage, while the blue area indicates that the non-minimal curvature interaction suppresses the quartic barrier. The light-green area indicates absolute stability for $m_{t}\leq\bar{m}_{t}$, where only the low-scale minimum is present. The metastable scenario involving two minima is realised in the white area for the general case of the Standard-Model only effective potential. Including the sextic interaction term $(H^{\dagger}H)^{3}/\Lambda^{2}$ leads to a shift in the absolute stability bound, indicated by the dotted lines for different values of the cut-off scale $\Lambda$. The model parameters have been fixed to $m_{h}=125.2\textrm{ GeV}$ and $\nu=20$, leading to an instability scale of $\mu_{\rm inst}=8.3\times 10^{13}\textrm{ GeV}$ and an effective-field-theory cut-off scale $\Lambda_{\xi}=9.2\times 10^{15}\textrm{ GeV}$ (dashed line).
@@
From a bottom-up perspective, such new physics would manifest itself as higher-dimensional operators suppressed by a cut-off scale $\Lambda\lesssim\Lambda_{\xi}$, determined by the thresholds and masses of the new particles as they are integrated out during the construction of the low-energy effective field theory. Interestingly enough, the inclusion of these non-renormalisable operators could potentially restore the convexity of the Higgs effective potential at high energies (see Figure 1), leading to the absolute stability of the electroweak vacuum irrespectively of the specific top-quark contribution. To investigate this possibility, we consider the impact of a prototypical non-renormalisable sextic operator, constructed from the gauge-invariant combination $H^{\dagger}H$ and the hypothetical cut-off scale of new physics $\Lambda$. The corresponding effective potential takes the form
$V_{\rm eff}(H)=(\xi R+m_{h}^{2})H^{\dagger}H+\lambda(H^{\dagger}H)^{2}+\frac{1}{\Lambda^{2}}(H^{\dagger}H)^{3}\;.$ (24)
@@
The inclusion of the sextic operator is expected to modify the $\beta$-functions of the theory Eichhorn:2015kea , requiring their rederivation for a fully rigorous treatment of the problem. However, since the threshold effects associated with non-renormalisable operators become significant only in the vicinity of the characteristic scale, we will proceed by retaining the original three-loop running in (3) while incorporating the sextic operator as a simple tree-level term, thereby neglecting its energy dependence Branchina:2013jra .
@@
Figure 7: Present-day SGWB generated by the non-minimally coupled Higgs in the presence of higher order operators in the effective potential. Different colours correspond to different choices of kination scale $\mathcal{H}_{\rm kin}$ in our benchmark simulations, while the cut-off scale of new physics has been set to $\Lambda=10^{15}\textrm{ GeV}$. The remaining model parameters have been set to $\nu=20$ and $m_{t}=171.3\textrm{ GeV}$. Dashed lines indicate the corresponding GW signals shown in Figure 3 for $m_{t}=171\textrm{ GeV}$ and $\mathcal{H}_{\rm kin}=10^{11},\,10^{12},\,10^{13}\textrm{ GeV}$.
@@
As illustrated in Figure 6, the presence of new physics at the scale $\Lambda$ shifts the Planckian minimum of the effective potential to smaller scales, or even eliminates it entirely. This figure provides a direct representation of the different phases of the model at the time of symmetry breaking, i.e. at the onset of kinetic domination. In the absence of new physics, the classical dynamics of the Higgs fields must be generally checked for instability within the parameter space corresponding to the white area, while absolute stability is ensured only for top quark masses below the critical value of $\bar{m}_{t}=171.035\textrm{ GeV}$. If the inflationary scale $\mathcal{H}_{\rm kin}$ is significantly higher than the instability scale $\mu_{\rm inst}$, the potential lacks a barrier and only the high-scale vacuum is present. The inclusion of sextic interactions stabilises the potential by restoring convexity at high scales. A large cut-off scale $\Lambda$ is enough to ensure stability for low top quark masses, while a smaller $\Lambda$ can stabilise the potential for a wider range of masses. As a result, the critical boundary at $\bar{m}_{t}$ shifts upwards, as indicated by the dashed lines in the figure, thus expanding the parameter space associated with absolutely stability. Notice that this analysis concerns only the overall vacuum structure of the potential at the time of symmetry breaking for a specific value of non-minimal coupling parameter and does not include the Higgs dynamics. [^7] In general, for the parameter space under consideration, the Hubble scale at the end of the first semi-oscillation of the field is about three orders of magnitude smaller than $\mathcal{H}_{\rm kin}$. Qualitatively speaking, this difference shifts the black boundary in Figure 6 to higher values of $\mathcal{H}_{\rm kin}$.
@@
Within the newly stabilised region, we turn once again to 3+1 classical lattice simulations. The main advantage of this numerical technique consists in the accurate description of the non-linear field dynamics and the prediction of the associated GW signal. The setup behind the implementation via the code $\mathcal{C}osmo\mathcal{L}attice$ Figueroa:2020rrl ; Figueroa:2021yhd is described in Appendix B. For this discussion, it suffices to state that the numerical model incorporates the complete potential in (24), along with its parametric three-loop running (3). The simulations are initialised at the time of symmetry breaking, i.e. at the transition between inflation and kination, and evolved for at least two e-folds on a kination background to capture the GW spectrum after the end of the non-linear regime. This implementation follows the established methods already tested in previous studies, including the SM Higgs scenario without new physics Laverda:2024qjt and the prototypical case of a scalar singlet at tree level Laverda:2023uqv ; Bettoni:2024ixe . We consider three benchmark points for $\mathcal{H}_{\rm kin}=10^{11},\,10^{12},\,10^{13}\textrm{ GeV}$, while we keep the scale of new physics at $\Lambda=10^{15}\textrm{ GeV}$. In order to assure the covering of all relevant momenta, we perform simulations on a large grid of $N=960$ lattice points per side.
The SGWB generated by the phase transition is presented in Figure 7. After extracting the spectrum at the timescale $z_{\rm rad}$ defined in (7), we rescale it to the present time using the formulas (16), assuming a typical value of $\lambda=10^{-3}$ for renormalisation scales in the range ${\mathcal{H}_{\rm kin}=10^{11}-10^{13}\textrm{ GeV}}$. For IR momenta, we recover the causal spectral slope proportional to $f^{3}$, while the UV portion of the spectrum is noticeably changed. The reason lies in the more efficient turbulent cascade Micha:2004bv towards UV modes induced by the sextic interaction. Consequently, the position of the peak is shifted to higher frequencies as compared to the absolutely-stable scenario (dashed lines in Figure 7) and its amplitude is reduced. Notice that this effect cannot be explained by the use of two different values of the top mass in the figure, since the peak frequency tends to decrease with higher top masses in the absolutely stable case, see Figure 6. The same turbulent process is responsible for the fast population of high-momentum modes, which flattens the overall tilt of the UV tail. These considerations are further supported by the gradual shift towards the UV of the BSM spectra with respect to the SM case as the scale of kination increases. Indeed, the SGWB matches well what one would expect from the absolute stability case for ${\mathcal{H}_{\rm kin}=10^{11}\textrm{ GeV}}$, since the classicalised Higgs fluctuations are well within the instability scale. However, at higher kination scales, the sextic operator becomes more and more relevant as the Higgs fluctuations become more and more classically unstable, thereby probing the new-physics scales. In essence, the presence of the additional higher-order operator rescues the parameter space excluded by stability constraints and leads to a GW signal that could not be produced within the SM effective potential. The three benchmark examples demonstrate that the SGWB signal possesses distinctive features at very high frequencies that provide us with useful information about the presence of non-renormalisable operators in the SM Higgs potential.
@@ 6 Conclusions
Vacuum stability is one of the most profound questions in fundamental physics, directly tied to the structure of the Higgs potential and the long-term fate of our Universe. One of the most significant outcomes of the Large Hadron Collider has been its ability to probe the stability of the electroweak vacuum by precisely measuring the Higgs boson and top quark masses, two key parameters that determine whether the vacuum is absolutely stable, metastable, or prone to decay over cosmological timescales. Intriguingly, current measurements place the SM vacuum near a critical boundary between stability and metastability, suggesting the possibility of new physics at higher energy scales. This delicate balance opens a compelling avenue for exploring BSM physics, as even slight modifications to the Higgs potential, such as new scalar fields, non-renormalisable interactions, or quantum gravity effects, could dramatically alter the fate of the vacuum.
The present work aims at probing the high-scale Higgs effective potential via GWs. The interplay of non-minimal gravitational couplings and a kination epoch results in a Hubble-induced phase transition that quickly reheats the Universe in a non-perturbative way while leaving a specific top-mass-dependent GW signature. We have shown that top quark masses compatible with absolute stability generate a strong signal, while higher top quark masses enforce more stringent upper bounds on the inflationary scale, thus weakening the overall signal. A simultaneous detection of the inflationary background by low-frequency experiments and of the peak of the phase transition’s spectrum by ultra-high-frequency detectors can quantify the model parameters. In this sense, the scenario under consideration offers a way to investigate gravitationally the magnitude of the top quark mass. Notably, while the peak frequency of the spectrum is strongly correlated with the top quark mass, the reheating temperature is predominantly determined by the effective tachyonic mass. Different cosmological constraints can act complementarily in characterising the gravitational and Standard-Model couplings of the Higgs at high scales.
A key aspect of the analysis is the extensive use of parametric formulas, derived from hundreds of classical lattice simulations, which enable us to predict the general shape of the SGWB as a function of the top quark mass, being this quantity the largest source of uncertainty in the stability of the electroweak vacuum. It has to be noted that, unlike in other works on GWs from vacuum instability Espinosa:2018eve , our setup does not rely on the temporary instability of the Higgs vacuum to generate a GW signal. On the contrary, we have shown that a GW signal is unavoidable even in the case of absolute stability because of the combination of non-minimal curvature couplings and a stiff post-inflationary expansion phase.
@@ Acknowledgments
We extend our gratitude to Dario Bettoni for engaging discussions that contributed to the preparation of this work. The numerical lattice simulations were carried out with the support of the Infraestrutura Nacional de Computação Distribuída (INCD) funded by the Fundação para a Ciência e a Tecnologia (FCT) and FEDER under the project 01/SAICT/2016 nº 022153. G. L. acknowledges support from a fellowship provided by “la Caixa” Foundation (ID 100010434) with fellowship code LCF/BQ/DI21/11860024, as well as the support of FCT through the grant with Ref. 2024.05847.BD. G. L. acknowledges also the financial support from FCT to the Center for Astrophysics and Gravitation-CENTRA, Instituto Superior Técnico, Universidade de Lisboa, through Project No. UIDB/00099/2020. J. R. is supported by a Ramón y Cajal contract of the Spanish Ministry of Science and Innovation with Ref. RYC2020-028870-I. This research was further supported by the project PID2022-139841NB-I00 of MICIU/AEI/10.13039/501100011033 and FEDER, UE.
Appendix A Summary of parametric formulas
In this Appendix, we present the full set of parametric formulas derived in Laverda:2023uqv ; Bettoni:2024ixe . In order to be as general as possible, we keep the notation in those works, namely the dependence on $\lambda$ and $\nu$ taken as tree-level parameters. At backreaction time $z_{\rm br}$, the amplitude and the energy density are given by

### v2

@@ 6 Stability beyond the Standard Model
As shown in Appendix A, the non-minimal coupling of the Higgs field to gravity makes the model under consideration intrinsically non-renormalisable, requiring it to be treated as an effective field theory valid up to a specific cut-off scale, $\Lambda_{\xi}=M_{P}/\xi$ Barbon:2009ya ; Burgess:2009ea ; Hertzberg:2010dc ; Burgess:2010zq ; Bezrukov:2010jz ; Bezrukov:2012hx ; Ren:2014sya ; Bezrukov:2014ipa ; Fumagalli:2017cdo ; Antoniadis:2021axu ; Mikura:2021clt ; Ito:2021ssc ; Escriva:2016cwl ; Karananas:2022byw , which may either signal the transition to a strongly coupled regime Aydemir:2012nz ; Calmet:2013hia ; Escriva:2016cwl ; Saltas:2015vsc or the emergence of new degrees of freedom beyond the SM content Giudice:2010ka (for a review in the context of Higgs inflation see Rubio:2018ogq ). For the non-minimal couplings considered in this work, $\Lambda_{\xi}$ is estimated to range from approximately between $10^{16}$ to $10^{17}$ GeV, i.e. typically well above the SM vacuum instability scale. However, the validity of the low-energy effective field theory could be further reduced in the presence of additional sub-Planckian degrees of freedom beyond the SM content.
@@
Figure 6: Impact of non-renormalisable sextic operators in the vacuum structure of the Higgs effective potential. The blue area indicates the parameter space where the non-minimal curvature interaction suppresses the quartic barrier. The light-green area indicates absolute stability for $m_{t}\leq\bar{m}_{t}$, where only the low-scale minimum is present. The metastable scenario involving two minima is realised in the white area for the general case of the Standard-Model only effective potential. Including the sextic interaction term $(H^{\dagger}H)^{3}/\Lambda^{2}$ leads to a shift in the absolute stability bound, indicated by the dotted lines for different values of the cut-off scale $\Lambda$. The model parameters have been fixed to $m_{h}=125.2\textrm{ GeV}$ and $\nu=20$, leading to an effective-field-theory cut-off scale $\Lambda_{\xi}=9.2\times 10^{15}\textrm{ GeV}$ (dashed line).
@@
From a bottom-up perspective, such new physics would manifest itself as higher-dimensional operators suppressed by a cut-off scale $\Lambda\lesssim\Lambda_{\xi}$, determined by the thresholds and masses of the new particles that are integrated out during the construction of the low-energy effective field theory. Interestingly enough, the inclusion of these so-far omitted non-renormalisable operators could potentially restore the convexity of the Higgs effective potential at high energies (see Figure 1), leading to the absolute stability of the electroweak vacuum irrespectively of the specific top-quark contribution. To investigate this possibility, we consider the impact of a prototypical non-renormalisable sextic operator, constructed from the gauge-invariant combination $H^{\dagger}H$ and the hypothetical cut-off scale of new physics $\Lambda$ as free parameter. The corresponding effective potential takes the form
$V_{\rm eff}(H)=\xi RH^{\dagger}H+\lambda(H^{\dagger}H)^{2}+\frac{1}{\Lambda^{2}}(H^{\dagger}H)^{3}\;.$ (31)
@@
Note that the inclusion of the sextic operator is expected to modify the $\beta$-functions of the theory Eichhorn:2015kea , requiring their rederivation for a fully rigorous treatment of the problem. However, since the threshold effects associated with non-renormalisable operators become significant only in the vicinity of the characteristic scale, we will proceed by retaining the original three-loop running in (10) while incorporating the sextic operator as a simple tree-level term, thereby neglecting its energy dependence Branchina:2013jra .
@@
Figure 7: GW signal after the end of the production phase in the scenario involving a non-minimally coupled Higgs with sextic self-interactions in the effective potential. Different colours correspond to different choices of the kination scale $\mathcal{H}_{\rm kin}$ in our benchmark simulations, while the cut-off scale of new physics has been set to $\Lambda=10^{15}\textrm{ GeV}$. The remaining model parameters have been set to $\nu=20$ and $m_{t}=171.3\textrm{ GeV}$. Dashed lines indicate the maximum and minimum GW signal shown in Figure 3 with $m_{t}=170.8\textrm{ GeV}$ and $\mathcal{H}_{\rm kin}=10^{11.5},\,10^{12}\textrm{ GeV}$. Momenta are normalised with respect to $\mathcal{H}_{\rm kin}$ and the simulations are well within the effective-field theory cut-off, namely $\langle h^{2}\rangle<10^{-2}M_{P}^{2}/\xi^{2}$.
@@
As illustrated in Figure 6, the presence of new physics at the scale $\Lambda$ shifts the super-Planckian minimum of the effective potential to smaller scales, or even eliminates it entirely. This figure provides a direct representation of the different phases of the model at the time of symmetry breaking, i.e. at the onset of kinetic domination. In the absence of new physics, the classical dynamics of the Higgs fields must be generally checked for instability within the parameter space corresponding to the white area, while absolute stability is ensured only for top-quark masses below the critical value of $\bar{m}_{t}=170.886\textrm{ GeV}$. If the inflationary scale $\mathcal{H}_{\rm kin}$ is sufficiently close to the instability scale $\mu_{\rm inst}$, the potential lacks a barrier due to the large contribution from the non-minimal coupling and only one high-scale vacuum is present. The inclusion of sextic interactions stabilises the potential by restoring convexity at sub-Planckian scales. When a barrier is present, a large cut-off scale $\Lambda$ is enough to ensure stability for low top-quark masses, while a smaller $\Lambda$ can stabilise the potential for a wider range of masses. As a result, the critical boundary at $\bar{m}_{t}$ shifts upwards, as indicated by the dashed lines in the figure, thus expanding the parameter space associated with absolute stability. When the barrier is fully suppressed, the higher-order operator provides the presence of a high-scale vacuum below the Planck mass. Notice that this analysis concerns only the overall vacuum structure of the potential at the time of symmetry breaking for a specific value of non-minimal coupling parameter and does not include the Higgs dynamics. [^9] In general, for the parameter space under consideration, the Hubble scale at the end of the first semi-oscillation of the field is about one order of magnitude smaller than $\mathcal{H}_{\rm kin}$. Qualitatively speaking, this difference shifts the black boundary in Figure 6 to higher values of $\mathcal{H}_{\rm kin}$.
@@
Within the newly stabilised region, we turn once again to 3+1 classical lattice simulations. The main advantage of this numerical technique consists in the accurate description of the non-linear field dynamics and the prediction of the associated GW signal. The setup behind the implementation via the code $\mathcal{C}osmo\mathcal{L}attice$ Figueroa:2020rrl ; Figueroa:2021yhd is described in Appendix C. For this discussion, it suffices to state that the numerical model incorporates the complete potential in (31), along with its parametric three-loop running (10). The simulations are initialised at the time of symmetry breaking, i.e. at the transition between inflation and kination, and evolved for at least two e-folds on a kination background to capture the GW spectrum after the end of the non-linear regime. This implementation follows the established methods already tested in previous studies, including the SM Higgs scenario without new physics Laverda:2024qjt and the prototypical case of a scalar singlet at tree level Laverda:2023uqv ; Bettoni:2024ixe . We consider four benchmark points for $\mathcal{H}_{\rm kin}\in[10^{11},\,10^{12.5}\textrm{ GeV}]$, while we keep the scale of new physics at $\Lambda=10^{15}\textrm{ GeV}$. In order to assure the covering of all relevant momenta, we perform simulations on a large grid of $N=960$ lattice points per side.
The GW signal generated by the phase transition is presented in Figure 7. The spectra are measured after the end of the large-oscillation non-linear phase that follows the tachyonic instability. Indeed, as was seen in Bettoni:2024ixe , the GW production is concentrated in this period, since the gradient energy density is at its peak. For IR momenta, we recover the causal spectral slope proportional to $f^{3}$. However, the frequency range of metric perturbations is shifted to higher values with respect to the GW signal shown in Figure 3. Interestingly, the position of the peak evolves from $\kappa\sim 120$ for $\mathcal{H}_{\rm kin}=10^{11}\textrm{ GeV}$ to $\kappa\sim 150$ for $\mathcal{H}_{\rm kin}=10^{12.5}\textrm{ GeV}$. This characteristic evolution with increasing inflationary scales is in contrast with the one expected in the pure SM scenario, where higher inflationary scales lead generally to smaller self-couplings and therefore lower peak frequencies. The difference is explained by the more efficient turbulent cascade Micha:2004bv towards UV modes induced by the sextic interaction, which quickly breaks down the large-amplitude fluctuations and in so doing it excites high-frequency modes more intensely than in the SM scenario. Moreover, one can also notice different slopes around the peak, which can constitute a telltale sign of the presence of new physics Dufaux:2010cf . Indeed, for high inflationary scales, the sextic operator becomes essential in guaranteeing the stability of the SM Higgs field and the stochastic system reflects the presence of such additional interactions. These bespoke features can shed light on the presence of non-renormalisable operators in the Higgs potential via a possible GW detection at very high frequencies.
@@ 7 Conclusions
Vacuum stability is one of the most profound questions in fundamental physics, directly tied to the structure of the Higgs potential and the long-term fate of our Universe. One of the most significant outcomes of the Large Hadron Collider has been its ability to probe the stability of the electroweak vacuum by precisely measuring the Higgs boson and top-quark masses, two key parameters that determine whether the vacuum is absolutely stable, metastable, or prone to decay over cosmological timescales. Intriguingly, current measurements place the SM vacuum near a critical boundary between stability and metastability, suggesting the possibility of new physics at higher energy scales. This delicate balance opens a compelling avenue for exploring BSM physics, as even slight modifications to the Higgs potential, such as new scalar fields, non-renormalisable interactions, or quantum gravity effects, could dramatically alter the fate of the vacuum.
The present work aims at probing the high-scale Higgs effective potential via GWs. The interplay of non-minimal gravitational couplings and a kination epoch results in a Hubble-induced phase transition that quickly reheats the Universe in a non-perturbative way while leaving a specific top-mass-dependent GW signature. We have shown that top-quark masses compatible with absolute stability generate a strong signal, while higher top-quark masses enforce more stringent upper bounds on the inflationary scale, thus weakening the overall signal. A simultaneous detection of the inflationary background at low-frequencies and of the peak of the phase transition’s spectrum at high frequencies can quantify the model parameters. In this sense, the scenario under consideration offers a way to investigate gravitationally the magnitude of the top-quark mass. Notably, while the peak frequency of the spectrum is strongly correlated with the top-quark mass, the (re)heating temperature is predominantly determined by the effective tachyonic mass. Different cosmological constraints can act complementarily in characterising the gravitational and Standard-Model couplings of the Higgs at high scales.
A key aspect of the analysis is the extensive use of parametric formulas, derived from hundreds of 3+1 classical lattice simulations, which enable us to predict the general shape of the SGWB as a function of the top-quark mass, being this quantity the largest source of uncertainty in the stability of the electroweak vacuum. It should be noted that, unlike other works on GWs from vacuum instability Espinosa:2018eve , our setup does not rely on the temporary instability of the Higgs vacuum to generate a GW signal. On the contrary, we have shown that a GW signal is unavoidable even in the case of absolute stability because of the combination of non-minimal curvature couplings and a stiff post-inflationary expansion phase.
@@ Acknowledgments
We extend our gratitude to Dario Bettoni for engaging discussions that contributed to the preparation of this work. We also thank Sebastian Ellis for the valuable comments on the first version of this manuscript. The numerical 3+1 classical lattice simulations were carried out with the support of the Infraestrutura Nacional de Computação Distribuída (INCD) funded by the Fundação para a Ciência e a Tecnologia (FCT) and FEDER under the project 01/SAICT/2016 nº 022153. G. L. acknowledges support from a fellowship provided by “la Caixa” Foundation (ID 100010434) with fellowship code LCF/BQ/DI21/11860024, as well as the support of FCT through the grant with Ref. 2024.05847.BD. G. L. acknowledges also the financial support from FCT to the Center for Astrophysics and Gravitation-CENTRA, Instituto Superior Técnico, Universidade de Lisboa, through Project No. UIDB/00099/2020. J. R. is supported by a Ramón y Cajal contract of the Spanish Ministry of Science and Innovation with Ref. RYC2020-028870-I. This research was further supported by the project PID2022-139841NB-I00 of MICIU/AEI/10.13039/501100011033 and FEDER, UE. Project carried out with the 2025 Leonardo Grant for Scientific Research and Cultural Creation from the BBVA Foundation. The BBVA Foundation is not responsible for the opinions, comments, and content included in the project and/or its resulting outcomes, which are the sole and exclusive responsibility of the authors
Appendix A Einstein-frame formulation and the EFT cut-off
The value of the cut-off scale entering the effective potential (31) is a priori unknown and depends on the various thresholds and masses of potential BSM particles that were integrated out to obtain the low-energy effective field theory. The non-renormalisability of the SM non-minimally coupled to gravity sets, however, an upper bound on this scale for a given value of $\xi$ Barbon:2009ya ; Burgess:2009ea ; Hertzberg:2010dc ; Burgess:2010zq ; Bezrukov:2010jz ; Bezrukov:2012hx ; Ren:2014sya ; Bezrukov:2014ipa ; Fumagalli:2017cdo ; Antoniadis:2021axu ; Mikura:2021clt ; Ito:2021ssc ; Escriva:2016cwl ; Karananas:2022byw . To see this explicitly, let us consider the action for the inflaton field $\phi$ and the subdominant Higgs field in the unitary gauge $H=(0,\,h/\sqrt{2})^{T}$,
$S=\int d^{4}x\sqrt{-g}\left[\left(\frac{M^{2}_{P}}{2}-\frac{1}{2}\xi h^{2}\right)R-\frac{1}{2}\partial^{\mu}h\partial_{\mu}h-\frac{\lambda}{4}h^{4}+\mathcal{L}_{\rm\phi}\right]\,.$ (32)
For the sake of concreteness and without any lack of generality, we will assume the inflaton field to display a standard Lagrangian density
$\mathcal{L}_{\phi}=-\frac{1}{2}\partial^{\mu}\phi\partial_{\mu}\phi-V(\phi)\,$, (33)
with $V(\phi)$ a runaway or quintessential potential able to support a sustained-enough period of kinetic domination immediately after the end of inflation. Performing a Weyl rescaling of the metric $g_{\mu\nu}\to\Omega^{-2}g_{\mu\nu}$ with conformal factor $\Omega^{2}=1-{\xi}h^{2}/M^{2}_{P}$, we obtain the equivalent Einstein-frame action
$S=\int d^{4}x\sqrt{-g}\left[\frac{M_{P}^{2}}{2}{R}-\frac{1}{2}\left(\frac{1}{\Omega^{2}}+\frac{6\xi^{2}h^{2}}{M_{P}^{2}\Omega^{4}}\right)\partial^{\mu}h\partial_{\mu}h-\frac{\lambda}{4}\frac{h^{4}}{\Omega^{4}}+\frac{1}{\Omega^{4}}{\tilde{\cal L}}_{\phi}\right]\,$, (34)
with
$\tilde{\cal L}_{\phi}=-\frac{1}{2}\,\Omega^{2}\partial^{\mu}\phi\partial_{\mu}\phi-V(\phi)\,.$ (35)
For non-minimal couplings $\xi\gg 1$ and field values $u\equiv\sqrt{\xi}h/M_{P}\ll 1$, this non-linear expression can be written
$S\simeq\int d^{4}x\sqrt{-g}\Bigg[\frac{M_{P}^{2}}{2}R$ $+\mathcal{L}_{\phi}-\frac{1}{2}\left(1+\frac{6h^{2}}{\Lambda_{\xi}^{2}}\right)\partial^{\mu}h\partial_{\mu}h-\frac{\lambda}{4}h^{4}$
$+\frac{1}{2}\frac{\xi h^{2}}{M^{2}_{P}}\,T^{(\phi)}+{\cal O}(u^{2}\,{\cal L}_{h},u^{4}{\cal L}_{\phi})\Bigg]\,$, (36)
where we have identified the cut-off-scale of the theory from the leading-order correction ($n=1$) to the Higgs kinetic term at large $\xi$ values,
$\Lambda_{\xi}=\frac{M_{P}}{\xi}\,$, (37)
explicitly isolated the $n=1$ correction to the inflaton Lagrangian density ${\cal L}_{\phi}$ and introduced the energy-momentum tensor of the inflaton field and its trace,
$T^{(\phi)}_{\mu\nu}=\partial_{\mu}\phi\partial_{\nu}\phi-g_{\mu\nu}\left(\frac{1}{2}\partial^{\rho}\phi\partial_{\rho}\phi+V(\phi)\right)\,,\quad\quad T^{(\phi)}\equiv T^{(\phi)\,\mu}{}_{\mu}=-\left(\partial^{\mu}\phi\partial_{\mu}\phi+4V(\phi)\right)\,.$ (38)
Notice that a hierarchy
$T^{(\phi)}\gg\frac{\xi h^{2}}{M^{2}_{P}}\,T^{(\phi)}\gg\frac{\xi h^{2}}{M^{2}_{P}}\,\mathcal{L}^{(h)},\frac{\xi^{2}h^{4}}{M^{4}_{P}}\mathcal{L}^{(\phi)}$ (39)
is realised in the perturbative expansion, which inherently suppresses any direct coupling between the Higgs and the inflaton that might arise in the Einstein frame even when the same tree-level coupling is absent from the Jordan-frame Lagrangian. Although the fifth term on the right-hand side of (36) is divided by $M_{P}^{2}$, it is not suppressed relative to the Higgs Lagrangian density for $T^{(\phi)}\gg T_{\xi=0}^{(h)}$. For an homogeneous inflaton field $\phi(t)$ in a flat FLRW background, $T^{(\phi)}$ becomes
$T^{(\phi)}=-\rho_{\phi}+3p_{\phi}=-(1-3w_{\phi})\rho_{\phi}\,$, (40)
with $\rho_{\phi}$ and $p_{\phi}$ the energy density and pressure of the inflaton field in this regime and $w_{\phi}=p_{\phi}/\rho_{\phi}$ the corresponding equation-of-state parameter. The dynamics associated with the action (36) coincides therefore with that assumed in the main text, provided the effect of the higher-dimensional operator suppressed by the cut-off scale $\Lambda_{\xi}$ can be neglected. Restricting ourselves to field values $h\ll\Lambda_{\xi}$, we have
$S\simeq\int d^{4}x\sqrt{-g}\left[\frac{M^{2}_{P}}{2}R+\mathcal{L}_{\phi}-\frac{1}{2}{\partial_{\mu}}{h}\partial^{\mu}h-\frac{1}{2}\xi Rh^{2}-\frac{\lambda}{4}h^{4}\right]\,$, (41)
with $R\simeq-T^{\phi}/M_{P}^{2}\simeq 3(1-3w_{\phi}){\cal H}^{2}$.
Appendix B Summary of parametric formulas from lattice simulations
In this Appendix, we summarise the most relevant quantities that characterise the course of a Hubble-induced phase transition in the Higgs sector. In order to maintain the approach followed in previous works, we consider the tree-level Higgs potential discussed in Laverda:2023uqv , with fixed $\lambda$ and $\nu$. There exist two essential timescales in the evolution of the spectator field after symmetry breaking
  • The backreaction time $z_{\rm br}$ is defined as the time at which self-interaction effects from the quartic potential term become non-negligible and the initial tachyonic phase is brought to an end. We use the following implicit definition
[…]
    where the auxiliary field $Y=h/a\sqrt{6\xi}\mathcal{H}_{\rm kin}$ is a rescaled conformal version of the Higgs field in unitary gauge and $M^{2}(z)=(\nu^{2}-1/4)/(z+\nu)^{2}$. The previous identity quantifies the time at which the effective frequency of the Higgs fluctuations vanish, thereby coinciding with the condition $Y^{\prime\prime}(z_{\rm br})=0$. The definition of backreaction time can be easily estimated thanks to the linear analysis performed in Bettoni:2019dcw . Here we only quote the main result following the standard procedure of solving mode-by-mode the equation of motion, namely the expression for the evolution of the Higgs fluctuations
    $Y^{2}_{\rm rms}(z)=\left(\frac{2\nu-1}{8\pi\nu}\right)^{2}\left(1+\frac{z}{\nu}\right)^{2\nu+1}\kappa_{*}^{2}(z)\,$, (43)
    with $\kappa_{*}=k/\sqrt{6\xi}\mathcal{H}_{\rm kin}=2\sqrt{\nu+1}\,/(2(z+\nu))$ a typical momentum scale. Thanks to a large number of numerical 3+1 classical lattice simulations, it has been possible to characterise the amplitude and energy density of the Higgs field at the backreaction timescale in terms of parametric formulas: